# Decoding for Punctured Convolutional and Turbo Codes: A Deep Learning Solution for Protocols Compliance

### v1

@@ 4.1 AWGN Channels: Benchmarking and Precision
In this section, we compare the performance of the proposed LSTM-based CNE with traditional decoders in an AWGN channel using BPSK modulation. The information bit lengths varied from 120 to 960, and the code rates ranged from 1/3 to 5/6. The BER was then measured for each combination of information bit length and code rate. For each Signal-to-Noise Ratio point, $10^{5}$ code blocks were simulated to ensure statistical reliability.
The simulation results for convolutional codes are depicted in Figure 10. The proposed decoding method exhibits comparable performance to the conventional Viterbi algorithm when the information bit length is short, such as 120 bits. However, as the bit length increases, the proposed method significantly outperforms the Viterbi decoder, achieving State-Of-The-Art (SOTA) performance. Notably, at an information bit length of 960 and a BER of $10^{-4}$, the proposed algorithm achieves a performance gain of over 1.0 dB across all code rates. This performance advantage stems from the CNE’s ability to exploit bidirectional LLR information, capturing both forward and backward dependencies among LLRs. In contrast, the Viterbi algorithm, constrained by limited traceback depth to minimize hardware complexity, fails to achieve a global optimum.
Figure 10: Performance comparison of the proposed CNE decoder and the conventional Viterbi algorithm for convolutional codes. The results demonstrate significant performance gains with increasing information bit lengths.
Figures 11 and 12 present the Turbo code simulation results. In Figure 11, the performance of the CNE decoder with 3 iterations is nearly identical to that of the conventional BCJR decoder with 6 iterations, even when the BCJR algorithm employs full traceback. This demonstrates the efficiency of the proposed method in reducing computational complexity while maintaining high decoding accuracy.
Figure 11: BER performance of Turbo codes using the proposed CNE decoder with 3 iterations compared to the BCJR decoder with 6 iterations. Results indicate comparable performance with reduced computational complexity.
Furthermore, Figure 12 shows the performance comparison at an identical iteration count of 3. The CNE decoder consistently outperforms the BCJR decoder across all information bit lengths and code rates. This superiority is attributed to the ability of the LSTM-based CNE to effectively model temporal dependencies in the data and refine LLR estimates through its iterative architecture, resulting in enhanced decoding performance.
@@
Figure 12: BER performance of Turbo codes comparing the proposed CNE decoder and the BCJR decoder, both with 3 iterations. The CNE decoder consistently achieves superior performance across all information bit lengths and code rates.
@@ 4.2 Rayleigh Channels: Generalization and Robustness
In this section, we evaluate the performance of the proposed LSTM-based CNE in a $4\times 4$ MIMO Rayleigh fading channel characterized by 3-tap multipath coefficients. The simulation utilized 16-Quadrature Amplitude Modulation (16-QAM) with Gray-coded mapping for signal transmission. The evaluation covered a range of information bit lengths from 120 to 960 and code rates from 1/3 to 5/6, mirroring the configurations used in the AWGN scenario. For Turbo codes, the BCJR algorithm performs up to 6 iterations, while the proposed CNE decoder is fixed at 3 iterations. To assess the impact of channel estimation accuracy, we simulated the system’s performance under both perfect channel state information and least squares channel estimation. To ensure statistical robustness, $10^{5}$ code blocks were simulated for each SNR point.
Tables 2 and 3 present the required $E_{b}/N_{0}$ values to achieve target BERs of $10^{-3}$ and $10^{-4}$, respectively, for convolutional and Turbo codes in Rayleigh fading channels. In the tables, $\overline{A}$ indicates that when $E_{b}/N_{0}>A$, the decoding performance converges to a BER of 0 before reaching the target BER, and "..." signifies that even at $E_{b}/N_{0}=40$ dB, the target BER remains unattained. Smaller $E_{b}/N_{0}$ values reflect better performance, and the best results are highlighted in bold.
From the results in Table 2 and 3, it is evident that the proposed CNE decoder achieves a significant performance improvement over traditional decoding algorithms in Rayleigh fading channels. Notably, under LS channel estimation, the CNE decoder demonstrates a substantial advantage, outperforming traditional algorithms by more than 5 dB in many scenarios. Furthermore, the CNE decoder even surpasses traditional decoders with perfect channel state information.
@@
Table 3: Simulation results for Turbo codes in a $4\times 4$ Rayleigh fading channel with 16-QAM modulation
The simulation results under Rayleigh fading channels provide critical insights into the strengths of the proposed LSTM-based CNE decoder, particularly in terms of generalization and robustness. The proposed CNE decoder achieves consistent performance gains under both AWGN and Rayleigh fading channels. This demonstrates its ability to generalize across different channel conditions, effectively learning the underlying relationships between received LLRs and transmitted bits, irrespective of the noise or fading environment. Unlike traditional decoders that rely on predefined assumptions and fixed traceback mechanisms, the CNE decoder dynamically adapts to complex channel conditions. Its LSTM-based architecture efficiently models temporal dependencies and leverages bidirectional information flow, enabling it to handle multipath interference and Rayleigh fading robustly.
@@ 5 Analysis
In this section, we analyze the proposed LSTM-based CNE decoder in terms of computational complexity and decoding latency. These two metrics are critical for evaluating the feasibility and efficiency of decoding algorithms in practical communication systems.
5.1 Computational Complexity Analysis
@@
The computational complexity of the proposed LSTM-based CNE decoder, as detailed in Eq. (13), arises from four key components: the input projection and gating operations, the batch normalization, the LSTM-based sequence processing, and the output projection.
@@
The first step involves projecting the input features $\boldsymbol{x}\in\mathbb{R}^{K\times D_{\text{in}}}$ through a fully connected layer with weight matrix $\boldsymbol{W}_{\text{proj}}\in\mathbb{R}^{D_{\text{embed}}\times D_{\text{in}}}$. Simultaneously, a gating mechanism modulates the features using the puncturing input $\boldsymbol{p}\in\mathbb{R}^{K\times D_{\text{in}}}$, with the gating weights parameterized by $\boldsymbol{W}_{\text{punc}}\in\mathbb{R}^{D_{\text{embed}}\times D_{\text{in}}}$. The complexity of this step, which involves matrix multiplications and sigmoid activations, is $\mathcal{O}\big{(}2K\cdot D_{\text{embed}}\cdot D_{\text{in}}+K\cdot D_{\text{embed}}\big{)}$.
Next, batch normalization is applied to the modulated features, which results in an output with dimension $\mathbb{R}^{K\times D_{\text{embed}}}$. Since batch normalization operates linearly over the embedding dimension $D_{\text{embed}}$, the computational complexity of this step is $\mathcal{O}(K\cdot D_{\text{embed}})$.
The batch-normalized features are then processed by the LSTM network, the core computational unit of the decoder. For a sequence length $K$, input size $D_{\text{embed}}$, and hidden state size $D_{\text{hidden}}$, the computational complexity of the bidirectional LSTM, accounting for the three gate operations (forget, input, and output gates) and the candidate memory cell, as well as the nonlinear activation functions, is given by:
@@
This complexity arises from matrix multiplications for each gate and memory cell, along with the non-linear activations, with the factor of 8 accounting for the gates and candidate cell, and the factor of 14 for the activation operations. The bidirectional structure doubles this complexity.
Finally, the LSTM outputs, $\boldsymbol{h}\in\mathbb{R}^{K\times 2D_{\text{hidden}}}$, are projected to the output space using the weight matrix $\boldsymbol{W}_{\text{out}}\in\mathbb{R}^{2D_{\text{hidden}}\times 1}$, resulting in the final output $\boldsymbol{y}\in\mathbb{R}^{K\times 1}$. The computational complexity of this step is $\mathcal{O}(2K\cdot D_{\text{hidden}})$.
Combining the complexities of the input projection, gating mechanism, batch normalization, LSTM processing, and output projection, the total computational complexity of the proposed CNE is:
@@
The computational complexity of the Viterbi algorithm is determined by the number of states in the trellis and the number of transitions evaluated at each time step. For a convolutional code with constraint length $L$, the trellis contains $2^{L-1}$ states. At each time step, the algorithm computes the path metrics for all $2^{L-1}$ states, resulting in a complexity of:
@@
Similarly, the BCJR algorithm, which performs forward and backward recursions over the trellis, requires computations proportional to the number of states in the trellis for each time step. Since both forward and backward recursions process $2^{L-1}$ states, and the Turbo decoder involves two component codes with $N_{\text{iter}}$ iterations, the total computational complexity of the BCJR algorithm is:

### v3

@@ 4.1 AWGN Channels: Benchmarking and Precision
In this section, we compare the performance of the proposed LSTM-based CNE with traditional decoders in an AWGN channel using BPSK modulation. The information bolck lengths varied from 120 to 960, and the code rates ranged from 1/3 to 5/6. The BER was then measured for each combination of information block length and code rate. For each SNR point, $10^{5}$ code blocks were simulated to ensure statistical reliability.
Figure 11: Performance comparison of the proposed CNE decoder and the conventional Viterbi algorithm for convolutional codes. The results demonstrate significant performance gains with increasing information block lengths.
The simulation results for convolutional codes are depicted in Figure 11. The proposed decoding approach exhibits comparable performance to the conventional Viterbi algorithm when the information block length is short, such as 120 bits. However, as the block length increases, the proposed CNE decoder significantly outperforms the Viterbi decoder [28], achieving state-of-the-art (SOTA) performance. Notably, at an information block length of 960 and a BER of $10^{-4}$, the proposed CNE decoder achieves a performance gain of over 1.0 dB across all code rates. This performance advantage stems from the CNE’s ability to exploit bidirectional LLR information, capturing both forward and backward dependencies among LLRs. In contrast, the Viterbi algorithm, constrained by limited traceback depth to minimize hardware complexity, fails to achieve a global optimum.
Figure 12: BER performance of Turbo codes with the proposed CNE decoder (3 iterations) versus BCJR decoders (3 and 6 iterations without a scaling factor). Results indicate comparable performance with reduced computational complexity.
In Figure 12, the proposed CNE decoder, operating with 3 iterations, achieves performance comparable to the conventional BCJR decoder with 6 iterations. This highlights the CNE’s efficiency in significantly reducing computational complexity while preserving high decoding accuracy. Moreover, Figure 12 shows the performance comparison at an identical iteration count of 3. The CNE decoder consistently surpasses the BCJR decoder across all evaluated information block lengths and code rates. This superiority is attributed to the ability of the LSTM-based CNE to effectively model temporal dependencies in the data and refine LLR estimates through its iterative architecture, resulting in enhanced decoding performance.
@@ 4.2 Rayleigh Channels: Generalization and Robustness
In this section, we evaluate the performance of the proposed CNE decoder in a $4\times 4$ MIMO Rayleigh fading channel characterized by 3-tap multipath coefficients. The simulation utilized 16-quadrature amplitude modulation (16-QAM) with Gray-coded mapping for signal transmission. The evaluation covered a range of information block lengths from 120 to 960 and code rates from 1/3 to 5/6, mirroring the configurations used in the AWGN scenario. For Turbo codes, the BCJR algorithm performs up to 6 iterations, while the proposed CNE decoder is fixed at 3 iterations. To assess the impact of channel estimation accuracy, we simulated the system’s performance under both perfect channel state information and least squares channel estimation. To ensure statistical robustness, $10^{5}$ code blocks were simulated for each SNR point.
Tables 3 and 4 present the required $E_{b}/N_{0}$ values to achieve target BERs of $10^{-3}$ and $10^{-4}$, respectively, for convolutional and Turbo codes in Rayleigh fading channels. In the tables, $\overline{A}$ indicates that when $E_{b}/N_{0}>A$, the decoding performance converges to a BER of 0 before reaching the target BER, and “...” signifies that even at $E_{b}/N_{0}=40$ dB, the target BER remains unattained. Smaller $E_{b}/N_{0}$ values reflect better performance, and the best results are highlighted in bold.
From the results in Table 3 and 4, it is evident that the proposed CNE decoder achieves a significant performance improvement over traditional decoding algorithms in Rayleigh fading channels. Notably, under LS channel estimation, the CNE decoder demonstrates a substantial advantage, outperforming traditional algorithms by more than 5 dB in many scenarios. Furthermore, the CNE decoder even surpasses traditional decoders with perfect channel state information.
@@
Table 4: Simulation results for Turbo codes in a $4\times 4$ Rayleigh fading channel with 16-QAM modulation. BCJR decoder without a scaling factor
The simulation results under Rayleigh fading channels provide critical insights into the strengths of the proposed CNE decoder, particularly in terms of generalization and robustness. The proposed approach achieves consistent performance gains under both AWGN and Rayleigh fading channels. This demonstrates its ability to generalize across different channel conditions, effectively learning the underlying relationships between received LLRs and transmitted bits, irrespective of the noise or fading environment. Unlike traditional decoders that rely on predefined assumptions and fixed traceback mechanisms, the CNE decoder dynamically adapts to complex channel conditions. Its LSTM-based architecture efficiently models temporal dependencies and leverages bidirectional information flow, enabling it to handle multipath interference and Rayleigh fading robustly.
@@ 5 Analysis
In this section, we examine the proposed CNE decoder, exploring its generalization capabilities, puncturing-aware embedding, computational complexity, decoding latency, and strategies for mitigating complexity and latency in practice. These aspects are crucial for understanding the algorithm’s performance, adaptability, and practical applicability in real-world communication systems.
5.1 Generalization Capabilities
Following the methodology described in the DeepTurbo paper, we reproduced its implementation (available at https://github.com/TechYan1990/deep-turbo/) and trained and evaluated it using the simulation parameters specified therein, as shown in Table 5. To ensure a fair comparison, we configured the CNE architecture to match DeepTurbo’s parameters, including a non-shared two-layer bidirectional gated recurrent unit with a hidden dimension $D_{\text{hidden}}=100$, embedding dimension $D_{\text{embed}}=5$, and six decoding iterations $N_{\text{iter}}=6$. Under these conditions, the primary architectural difference between CNE and DeepTurbo lies in the embedding strategy: CNE projects the systematic bits ($\mathbf{llr}_{s}$) and both parity sequences ($\mathbf{llr}_{z}$, $\mathbf{llr}_{z^{\prime}}$) into a high-dimensional space, whereas DeepTurbo only projects posterior information.
Table 5: Parameters of DeepTurbo and CNE simulations to analyze generalization capabilities
DeepTurbo CNE Turbo RNN Architecture Non-Shared 2-Layer Bi-GRU Non-Shared 2-Layer Bi-GRU Training Epochs 1000 1000 Batch Size 128 128 Batches per Epoch 128 128 Training Information Block Length 100 100 Inference Information Block Length 1000 1000 Number of Inference Code Blocks 10000 10000 Learning Rate 0.001 0.001 Optimizer Adam Adam Loss Function BCE BCE Training SNR 0 dB 0 dB Training Code Rate $1/3$ $1/3$ Posterior Feature Size $F_{\text{s}}$ 5 N/A Embedding Dimension $D_{\text{embed}}$ N/A 5 Hidden Dimension $D_{\text{hidden}}$ 100 100 Number of Iterations $N_{\text{iter}}$ 6 6 Puncturing-Aware Embedding N/A None
Figure 13: BER performance of Turbo codes decoded by CNE, DeepTurbo, and BCJR (6 iterations without a scaling factor) on an AWGN channel with information block length 1000.
Both CNE and DeepTurbo were trained on an AWGN channel with a information block length of 100 and evaluated at a block length of 1000 to assess generalization. The BCJR algorithm, used as a baseline, was configured with six iterations. The results, presented in Figure 13, demonstrate that DeepTurbo exhibits poor generalization at high SNRs, manifesting as a BER floor, whereas CNE maintains robust performance without this limitation.
These findings indicate that CNE’s superior generalization to different code lengths is not merely a result of increased parameters, as both architectures use identical parameter counts in this experiment. Instead, the key preprocessing step contributing to generalization is CNE’s comprehensive embedding of $\mathbf{llr}_{s}$, $\mathbf{llr}_{z}$, and $\mathbf{llr}_{z^{\prime}}$ into a high-dimensional space. This approach enhances the neural network’s expressive power, enabling it to capture complex relationships within the codeword more effectively than DeepTurbo’s posterior-only embedding.
5.2 Puncturing-Aware Embedding
Table 6: Parameters of DeepTurbo and CNE simulations to analyze puncturing-aware embedding
DeepTurbo CNE Turbo RNN Architecture Non-Shared 2-Layer Bi-GRU Shared 2-Layer Bi-LSTM Training Epochs 1000 1000 Batch Size 128 128 Batches per Epoch 128 128 Training Information Block Length 120 120 Inference Information Block Length 120 120 Number of Inference Code Blocks 10000 10000 Learning Rate 0.001 0.001 Optimizer Adam Adam Loss Function BCE BCE Training SNR 0 dB 0 dB Training Code Rate $1/3$ $1/3$ Fine-Tuning Code Rate $1/3,1/2,2/3,3/4$ $1/3,1/2,2/3,3/4$ Posterior Feature Size $F_{\text{s}}$ 5 N/A Embedding Dimension $D_{\text{embed}}$ N/A 64 Hidden Dimension $D_{\text{hidden}}$ 256 256 Number of Iterations $N_{\text{iter}}$ 6 3 Puncturing-Aware Embedding N/A None
Figure 14: BER performance of Turbo codes decoded by CNE (without the puncturing-aware embedding module), and BCJR (6 iterations without a scaling factor) on an AWGN channel with information block length 120.
Figure 15: BER performance of Turbo codes decoded by DeepTurbo, and BCJR (6 iterations without a scaling factor) on an AWGN channel with information block length 120.
This secion we persent an ablation study on the effect of the puncturing-aware embedding module and the gain of the proposed CNE over a DeepTurbo-like decoder for punctured codes, providing a detailed comparison of decoding performance in terms of BER. This study evaluates CNE (without the puncturing-aware embedding module) against a DeepTurbo-like decoder and the traditional BCJR algorithm under identical training conditions.
For a fair comparison, both CNE and DeepTurbo are configured with their respective optimal parameters, where the hyperparameters of DeepTurbo are set according to [29], and the hidden dimension of DeepTurbo is increased from 100 in the original paper to 256 for a fair comparison, as summarized in Table 6. During fine-tuning, code rates of 1/3, 1/2, 2/3, and 3/4 are used, with the 5/6 code rate reserved as an unseen rate to assess generalization. Training and inference are conducted in an AWGN channel.
  1. 1.
    CNE vs. BCJR: Figure 14 compares the BER performance of CNE (without the puncturing-aware embedding module) against the BCJR algorithm. When the puncturing-aware embedding module is disabled, CNE exhibits limited generalization to the unseen 5/6 code rate, with a performance degradation of approximately 1.5 dB at BER = $10^{-4}$. For other code rates (1/3, 1/2, 2/3, and 3/4), the BER performance of CNE without the module is comparable to the BCJR algorithm, indicating that the absence of puncturing awareness hampers its ability to adapt to unseen rates.
  2. 2.
    DeepTurbo vs. BCJR: Figure 15 shows the BER performance of the DeepTurbo decoder compared to the BCJR algorithm. DeepTurbo exhibits minimal performance loss for the 1/3 code rate but suffers from a BER floor, indicating limited error correction capability at higher SNRs. For other code rates (1/2, 2/3, 3/4, and 5/6), DeepTurbo shows significant performance degradation and fails to demonstrate meaningful generalization, underscoring its inability to handle punctured codes effectively.
  3. 3.
    Key Insight: From the simulation results in Figure 14, it is evident that, although the absence of the puncturing-aware embedding module severely degrades the performance of CNE for the 5/6 code rate, it still exhibits some generalization capability. In contrast, the simulation results in Figure 15 show that DeepTurbo, even after fine-tuning, demonstrates no generalization to the 5/6 code rate. This indicates that the generalization ability for code rates is not solely due to the puncturing-aware embedding module but also stems from the contribution of projecting the systematic bits ($\mathbf{llr}_{s}$) and parity sequences ($\mathbf{llr}_{z}$, $\mathbf{llr}_{z^{\prime}}$) into a high-dimensional embedding space.
The ablation study confirms that the puncturing-aware embedding module is a pivotal component of the proposed CNE, enabling superior generalization to unseen code rates (e.g., 5/6) compared to both CNE without the module and a DeepTurbo-like decoder. The module’s ability to encode puncturing patterns into the latent space, combined with high-dimensional embedding of systematic and parity sequences, drives CNE’s robustness and protocol compatibility.
Figure 16: BER performance of Turbo codes decoded by CNE, DeepTurbo, and BCJR (6 iterations without a scaling factor) on an AWGN channel with information block lengths of 120, 240, 480, and 960 at a fixed non-punctured code rate of 1/3.
Based on the analysis above, due to DeepTurbo’s limited generalization ability across information block lengths and rates, we trained separate DeepTurbo models for each information block length ($K=120,240,480,960$) with a fixed non-punctured code rate of 1/3. The hyperparameters were configured according to Table 6, with the fine-tuning step omitted. The simulation results, shown in Figure 16, indicate that DeepTurbo exhibits a performance gap of approximately 0.2 dB compared to CNE at a BER of $10^{-4}$ under fixed code length and rate conditions.
5.3 Computational Complexity
@@
The computational complexity of the proposed CNE decoder, as detailed in Eq. (13), arises from four key components: the input projection and gating operations, the batch normalization, the LSTM-based sequence processing, and the output projection.
@@
The input features $\bm{x}\in\mathbb{R}^{K\times D_{\text{in}}}$ are projected using $\bm{W}_{\text{proj}}\in\mathbb{R}^{D_{\text{embed}}\times D_{\text{in}}}$, while puncturing patterns $\bm{p}\in\mathbb{R}^{K\times D_{\text{in}}}$ are processed with $\bm{W}_{\text{punc}}\in\mathbb{R}^{D_{\text{embed}}\times D_{\text{in}}}$. This step, involving matrix multiplications and sigmoid activations, incurs a complexity of $\mathcal{O}(2K\cdot D_{\text{embed}}\cdot D_{\text{in}}+K\cdot D_{\text{embed}})$.
Following this, batch normalization is applied to the modulated features in $\mathbb{R}^{K\times D_{\text{embed}}}$, with a complexity of $\mathcal{O}(K\cdot D_{\text{embed}})$. The core computational unit, a bidirectional LSTM, processes sequences of length $K$ with input size $D_{\text{embed}}$ and hidden state size $D_{\text{hidden}}$, involving three gates and a candidate memory cell. Its complexity, dominated by quadratic scaling with $D_{\text{hidden}}$, is:
@@
Finally, the LSTM output $\bm{h}\in\mathbb{R}^{K\times 2D_{\text{hidden}}}$ is projected using $\bm{W}_{\text{out}}\in\mathbb{R}^{2D_{\text{hidden}}\times 1}$, with a complexity of $\mathcal{O}(2K\cdot D_{\text{hidden}})$. The total CNE complexity is:
@@
For comparison, the Viterbi algorithm for a convolutional code with constraint length $L$ has a complexity of:
@@
The BCJR algorithm for Turbo decoding, with $N_{\text{iter}}$ iterations, has a complexity of:
@@
Table 7 compares the complexity of CNE and DeepTurbo decoders, focusing on trainable parameters and multiply-accumulate operations (MACs) per decoded bit, evaluated using Thop [41], a tool designed to measure the MACs of neural networks. Hyperparameters are listed in Table 6, except that CNE includes the puncturing-aware embedding module.
Table 7: Complexity comparison of CNE and DeepTurbo decoder
Number of Iterations Trainable Parameter MACs/decoded bit CNE Convolutional N/A 2,237,441 2,245,632 CNE Turbo 3 6,715,398 16,168,550 DeepTurbo 3 9,554,016 11,516,463 DeepTurbo 6 19,108,026 23,032,921
The CNE Turbo decoder, with shared weights between CNE0 and CNE1, significantly reduces trainable parameters compared to DeepTurbo’s non-shared SISO weights. For 3 iterations, DeepTurbo requires approximately 1.42 times the parameters of CNE Turbo, and for 6 iterations, this increases to about 2.85 times. The shared-weight design of CNE enhances efficiency by lowering storage and computational overhead.
In terms of computational complexity, DeepTurbo at 3 iterations requires roughly 0.71 times the MACs per decoded bit of CNE Turbo, due to its lower projection dimension ($F_{\text{s}}=5$ vs. CNE’s $D_{\text{embed}}=64$). However, at 6 iterations—DeepTurbo’s optimal configuration—it demands approximately 1.42 times the MACs per decoded bit of CNE Turbo. This demonstrates that CNE achieves superior performance with competitive complexity using fewer iterations.
5.4 Decoding Latency
Decoding latency is critical for real-time communication systems, and the CNE decoder’s latency stems from input projection and gating, batch normalization, LSTM processing, and output projection.
The input projection maps $\bm{x}\in\mathbb{R}^{K\times D_{\text{in}}}$ with $\bm{W}_{\text{proj}}\in\mathbb{R}^{D_{\text{embed}}\times D_{\text{in}}}$, while modulating with $\bm{p}\in\mathbb{R}^{K\times D_{\text{in}}}$ via $\bm{W}_{\text{punc}}\in\mathbb{R}^{D_{\text{embed}}\times D_{\text{in}}}$, using parallel matrix-vector multiplications with latency:
$T_{\text{proj}}=t_{\text{mat}}(D_{\text{embed}},D_{\text{in}})$ (28)
Batch normalization then normalizes features in $\mathbb{R}^{K\times D_{\text{embed}}}$ in parallel, with latency:
$T_{\text{BN}}=t_{\text{bn}}(D_{\text{embed}})$ (29)
The sequential LSTM, processing $K$ time steps, dominates latency due to its sequential nature:
The final output projection maps $\bm{h}\in\mathbb{R}^{K\times 2D_{\text{hidden}}}$ with $\bm{W}_{\text{out}}\in\mathbb{R}^{2D_{\text{hidden}}\times 1}$, with latency:
$T_{\text{out}}=t_{\text{mat}}(1,2D_{\text{hidden}})$ (31)
The total CNE latency is:
$T_{\text{CNE}}=T_{\text{proj}}+T_{\text{BN}}+T_{\text{LSTM}}+T_{\text{out}}$ (32)
The Viterbi algorithm’s latency, scaling with trellis states, is:
$T_{\text{Viterbi}}=K\cdot t_{\text{state}}(2^{L-1})$ (33)
The BCJR algorithm, with forward and backward recursions and $N_{\text{iter}}$ iterations, has latency:
$T_{\text{BCJR}}=2\cdot N_{\text{iter}}\cdot K\cdot t_{\text{state}}(2^{L-1})$ (34)
In addition, we evaluated the proposed CNE decoder using Torch-TensorRT [42], a tool designed to accelerate neural network inference on NVIDIA GPUs. The evaluation was performed with a floating-point precision of float32, based on 1000 trials with an information block length of 120. The results show that the proposed CNE convolutional code decoder has an average inference latency of $0.116\,\mu\text{s/decoded bit}$, while the Turbo code decoder exhibits an average inference latency of $0.867\,\mu\text{s/decoded bit}$.
5.5 Mitigating Complexity and Latency in Practice
While the proposed CNE decoder demonstrates superior performance, its computational complexity and latency, particularly due to the sequential nature of the LSTM, pose challenges for real-time and resource-constrained communication systems. To address these drawbacks and ensure practical deployment beyond reliance on GPU acceleration, a multi-faceted approach is presented, encompassing model optimization, hardware acceleration, and scheduling efficiency.
  1. 1.